Period of |sinπx| Nikhil Pandey Nikhil pandey jaisa hojay ggvnfjnvvfxvvxbbvxnbxnv

Grade:12th pass

1 Answers

Arun
6 years ago
we know that period of sinx is 2\pi
but if we take mod, graph of sinx will always be above y axis
hence peiod of |sinx| will be \pi.
but here period of |sin\pix| is asked then, its answer will be = \pi/\pi= 1